# Mobile communication, Wireless communication, Signal coverage, Data processing, Quality of Experience

### v1

@@ Abstract
Effective assessment of mobile network coverage and the precise identification of service weak spots are paramount for network operators striving to enhance user Quality of Experience (QoE). This paper presents a novel framework for mobile coverage and weak spot analysis utilising crowdsourced QoE data. The core of our methodology involves coverage analysis at the individual cell (antenna) level, subsequently aggregated to the site level, using empirical geolocation data. A key contribution of this research is the application of One-Class Support Vector Machine (OC-SVM) algorithm for calculating mobile network coverage. This approach models the decision hyperplane as the effective coverage contour, facilitating robust calculation of coverage areas for individual cells and entire sites. The same methodology is extended to analyse crowdsourced service loss reports, thereby identifying and quantifying geographically localised weak spots. Our findings demonstrate the efficacy of this novel framework in accurately mapping mobile coverage and, crucially, in highlighting granular areas of signal deficiency, particularly within complex urban environments.
@@ 1 Background
The empirical assessment of mobile network coverage predominantly relies on two distinct methodologies: traditional drive testing and contemporary crowdsourced data analysis. Drive testing, a long-established approach, involves systematic data collection utilising specialised Radio Frequency (RF) measurement equipment. As exemplified by regulatory bodies like the UK’s Office of Communications (Ofcom) Ofcom (2024a), this often includes sophisticated scanning receivers and calibrated devices capable of concurrently measuring key performance indicators across multiple mobile network operators and technologies (e.g., 2G, 3G, 4G, and 5G) along predefined routes Ofcom (2024b). This method affords a high degree of control over testing parameters and precise geolocation, enabling the acquisition of granular data essential for ensuring regulatory compliance and informing policy making. The primary strengths of drive testing lie in its precision, the depth of diagnostic information obtainable, and the ability to perform controlled, repeatable measurements. However, its operational costs are substantial due to specialised equipment and personnel, and its spatial-temporal coverage is inherently limited by logistical constraints, often restricting data to major transport corridors.
In contrast, crowdsourced data analysis leverages measurements collected from mobile apps on multitude of consumer smartphones. This approach offers unparalleled scalability, providing extensive geographic and temporal coverage at a significantly lower operational cost than drive testing. The data typically includes network coverage, internet connectivity, device information, etc. In particular, crowdsourced apps enable extensive sampling in indoor locations which are often inaccessible to conventional drive testing. Notably, crowdsourced apps even record instances of no signal, immediately logging a location when a phone has no coverage. These "no-service" data points highlight complete coverage holes that traditional methods might miss. Analysing these areas may reveal important socioeconmic patterns Koutroumpis and Leiponen (2016) to drive discussions on evidence-based policymaking processes.
@@
With crowdsourced data, we compare and contrast a geometric approach for signal coverage analysis and then propose a novel machine learning method. We also outline a time-series based validation strategy and examine the results. Finally, we consider practical engineering applications of improved coverage analysis and their business impact.
2 Methdology
@@
One traditional approach to estimating mobile coverage areas from crowdsourced data is to use computational geometry. For example, constructing a convex hull around all the geolocated points where service was observed. The convex hull provides a simple approximate boundary of the network’s coverage. If a geographic area lies outside this hull, it can be flagged as a potential coverage hole. This method is conceptually straightforward and computationally efficient. Yet, convex hull method tends to overgeneralise as it cannot represent concavities, internal holes, or irregular coverage boundaries, and it is strongly affected by outliers or widely spaced points Neidhardt et al. (2013); s23010352.
@@
In our approach, we model coverage estimation as a one‑class classification (novelty‑detection) problem. given geolocated measurements that evidence usable service (the "inliers"), learn the support of that distribution and treat points outside as likely weak or no coverage. The One‑class SVM (OC-SVM) algorithm Schölkopf et al. (2001) estimates a function $f(x)$ whose sign indicates membership in the learned support. The prediction $f(x)\geq 0$ denotes coverage area, where $f(x)<0$ denotes the otherwise. Unlike convex geometric baselines, the kernelised decision boundary can be highly non‑convex, capturing concavities and internal holes driven by terrain, clutter, or shadowing. In practice we use the Radial Basis Function (RBF) kernel to obtain smooth, locality‑aware boundaries that wrap around dense regions of positive evidence without being pulled by isolated outliers. This mirrors common use of one-class SVM in anomaly detection, where the algorithm learns the region occupied by "normal" data and flags deviations as anomalies.
@@
A key advantage of OC‑SVM is the soft boundary controlled by the hyperparameter $\nu\in(0,1]$. The $\nu$ parameter simultaneously sets an upper bound on the fraction of training errors (points allowed outside the learned support) and a lower bound on the fraction of support vectors. It controls the trade-off between overfitting (making the boundary too tight, excluding real coverage) and underfitting (making the boundary too loose, including false coverage). A small $\nu$ yields a tighter boundary that may exclude some true coverage points, while a large $\nu$ produces a looser boundary that may include false positives. Thus, $\nu$ governs the model’s sensitivity to outliers and the complexity of the learned support region.
@@
Figure 1: Mobile signal coverage area of a single cell tower. On the left is the convex hull approach (a), which overgeneralises and includes large areas of no coverage. On the right is the OC-SVM approach, which captures concavities and internal holes to better represent the true coverage area. The right figure (b) is a smoothed, non-convex boundary that tightly wraps around the empirical coverage points, calculated using OC-SVM with RBF kernel.
@@
Meanwhile, the RBF width is also governed by $\gamma$ in $K(x,x^{\prime})=\exp(-\gamma\lVert x-x^{\prime}\rVert^{2})$. A large $\gamma$ yields a narrow influence radius: each support vector only affects classification in its immediate vicinity, which allows the learned decision boundary to become sensitive and potentially overfitting to noise. A small $\gamma$ produces a broad influence radius, where each support vector affects a large region. This leading to a smoother, more general boundary that may underfit fine spatial detail. Small $\gamma$ produces a broad, smoother boundary that generalizes over larger spatial scales; large $\gamma$ yields a complex boundary that can overfit local points.
@@
To select the best hyperparameters, we tune $(\nu,\gamma)$ by temporal cross validation using held‑out time slices. We partition the data into training and validation sets along the temporal dimension (e.g., train on January and validate on Feburary). This simulates real-world deployment where we want to predict future coverage based on past data. For each candidate $(\nu,\gamma)$ pair, we train the OC‑SVM on the training set and evaluate its performance on the validation set using metrics like recall (true positive rate) and precision (positive predictive value). We select the hyperparameters that yield the best trade-off between recall and precision on the validation set, ensuring robust generalisation to unseen data.
@@
In our approach, we partition the crowdsourced measurement data by signal levels and train a separate OC‑SVM boundary for each level. The intuition is that the spatial support of good coverage points will differ (be more conservative) from that of weaker signal and so by modeling each level’s support separately, we can obtain layered (nested) boundaries that demarcate zones of stronger vs weaker coverage. Each OC‑SVM is responsible for estimating the region in which that signal level is reliably observed. In deployment, one can interpret a location’s highest predicted signal level by querying which OC‑SVM boundary it falls inside.
@@
We compare our OC‑SVM approach against the convex hull baseline by evaluating both methods on the same temporal validation splits. We assess their ability to correctly identify known coverage areas (true positives) and avoid falsely predicting coverage in known no-service areas (false positives). Metrics like recall, precision, and F1-score provide a quantitative comparison of their performance. We expect the OC‑SVM to outperform the convex hull by better capturing complex, non-convex coverage boundaries and reducing false positives due to its learned, data-driven nature.
@@ 3 Results
We conducted a grid search over $\nu\in\{0.02,0.04,0.06,0.08\}$ and $\gamma\in\{1\times 10^{4},2\times 10^{4},3\times 10^{4},4\times 10^{4}\}$ using time-slice cross-validation. As illustrated in Figure 2, lower values of $\nu$ produced tighter decision boundaries with fewer outliers, but often overfit the training data. In contrast, larger $\nu$ values allowed for a more points to lie outside the boundary.
Figure 2: Effect of varying hyparameters $(\nu,\gamma)$ on the shape and fidelity of the OC-SVM boundary. Note the trade-off between smoothness and fragmentation.
@@
We systematically evaluate the performance of OC-SVM boundaries using various hyperparameters against the convex hull baseline. The choice of F1 score is particularly useful in our context as it provides a single metric that balances both precision and recall. This is important because, in coverage area estimation, we want to minimize both false positives (areas incorrectly marked as covered) and false negatives (missed coverage areas). By using the F1 score, we can ensure that our model performs well across both metrics, rather than optimising for one at the expense of the other. The F1 score formula is given by:
@@
Catagory Signal Level (dBm) Convex Hull OC-SVM
@@
We performed cross-validation by partitioning the data into training and validation sets along the temporal dimension. For each candidate pair of hyperparameters $(\nu,\gamma)$, we train the OC-SVM on the training set and evaluate its performance on the validation set. The convex hull method is applied to the same training data for a fair baseline comparison. Each model is trained for individual cell towers to capture their unique coverage characteristics. We used 4,000+ cell towers across England for our analysis, ensuring a diverse representation of urban and rural environments. The results in Table 1 and Figure 3 are averaged over all cell towers. We observed that the OC-SVM approach consistently outperforms the convex hull approach, particularly in scenarios with variable or poor signal levels where coverage boundaries are more complex. This demonstrates the effectiveness of the OC-SVM in capturing non-convex coverage areas and reducing false positives.
@@ 4 Discussion
This study set out to address the challenge of accurately estimating mobile network coverage areas using crowdsourced data. The primary aim was to evaluate whether the OC-SVM could provide more precise and reliable coverage boundaries compared to a more traditional geometric method using convex hull. By doing so, we sought to improve the identification of coverage gaps and weak spots, ultimately supporting better network planning and QoE.
The results demonstrate that the OC-SVM approach consistently improves on the accuracy of coverage boundary estimation compared to the convex hull approach. This is particularly significant in urban environments where coverage gaps are often hidden by complex terrain and building structures. The kernelised decision boundary allows for capturing complex, non-convex shapes that more accurately reflect real-world coverage. This results in improved precision, as the model is less likely to include areas without service, and enhanced recall, as it can better identify true coverage areas.
There have been multiple attempts at utilising Machine Learning (ML) techniques to model and predict enhance traditional path loss models Zhang et al. (2019); Sousa et al. (2021); Isabona et al. (2022). However, the application of a OC-SVM classification method utilising crowdsourced data for coverage estimation is a novel contribution that extends the application of ML in wireless communications. This approach leverages the strengths of ML in handling complex, non-linear relationships in data, making it well-suited for capturing the intricacies of mobile network coverage that traditional methods may overlook. Introducing crowdsourced data into the modelling process also provides a more empirical basis for understanding real-world coverage patterns, as opposed to relying solely on theoretical models or limited test drive data that is both expensive and time-consuming to collect. Additionally, the utilisation of crowdsourced data allows for a more dynamic and up-to-date representation of network coverage, reflecting evolving changes in the environment and user behaviour that static models may miss, such as new building developments, changes in user density, and seasonal variations in environment.
@@
The improved accuracy of coverage boundary estimation has significant implications for network planning and optimisation. By accurately identifying coverage gaps and weak spots, operators can make informed decisions about where to deploy additional resources to address service gaps, such as new cell sites or signal boosters, and for considerations into appropriate spectrum bandings to apply in both rural and urban areas. This targeted approach can lead to more efficient use of resources and ultimately enhance the QoE for end users.
@@ 5 Conclusion and Future Work
This study proposes OC-SVM as a novel method for mobile network coverage analysis using crowdsourced geolocation data. We compared the proposed approach against a geometric convex hull baseline. We demonstrated that kernelised OC-SVM can effectively capture complex, non-convex coverage boundaries that better reflect real-world signal propagation, particularly in urban environments with significant clutter, terrain and shadowing effects. With appropriate tuning, the SVM boundaries achieve a balance between coverage inclusivity and the exclusion of spurious outliers, enabling the estimation of coverage boundaries at various signal strength levels. The results confirm that crowdsourced data, despite inherent heterogeneity and noise, can be transformed into reliable coverage models when combined with suitable machine learning techniques. This offers direct value both to operators seeking to prioritise investment and to regulators monitoring equitable service provision.
@@
Besides, kernelised OC-SVM natively supports multi-dimensional inputs, our proposed approach can be therefore extended to 3D geolocation data (ie. longitude, latitude, altitude). This could further enrich coverage analysis, particularly in dense urban environments where verticality plays a key role.
@@
In summary, the proposed method demonstrates that machine learning approach can be used to estimate mobile coverage boundaries from empirical crowdsourced data. This opens up new possibilities for network analysis, optimisation, and regulation. With further refinement and extension, this framework has a prospect of becoming a standard tool for mobile network coverage assessment in the era of machine learning.
@@
Each measurement record includes information such as timestamp, geographic coordinates, radio technologies (2G, 3G, 4G…), and key performance indicators including received signal strength, throughput, latency, and service availability. These metrics reflect the network’s performance as experienced by users, rather than inferred from network-side counters. All data are anonymised, aggregated, and processed in compliance with data protection and privacy frameworks before being made available for analysis.
For this study, a synthetically generated dataset based on UK NetPerform was used Feehily et al. (2025). The synthetic data comprise geolocated mobile signal measurements spanning across multiple months, capturing a broad range of environmental and temporal conditions. The synthetic dataset underwent standard pre-processing, including coordinate validation, outlier removal, and spatial filtering to exclude duplicate events.

### v2

@@ Abstract
Effective assessment of mobile network coverage and the precise identification of service weak spots are paramount for network operators striving to enhance user Quality of Experience (QoE). This paper presents a novel framework for mobile coverage and weak spot analysis utilising crowdsourced QoE data. The core of our methodology analyses coverage at the individual cell (antenna) level using empirical geolocation data before aggregating to the site level. A key contribution of this research is the application of the One-Class Support Vector Machine (OC-SVM) algorithm to calculate mobile network coverage. This approach models the decision hyperplane as the effective coverage contour, facilitating robust determination of coverage areas for individual cells and entire sites. The same methodology is extended to analyse crowdsourced service loss reports, thereby identifying and quantifying geographically localised weak spots. Our findings demonstrate the efficacy of this framework in accurately mapping mobile coverage and, crucially, in highlighting granular areas of signal deficiency, particularly within complex urban environments.
@@ 1 Background
The empirical assessment of mobile network coverage predominantly relies on two distinct methodologies: traditional drive testing and contemporary crowdsourced data analysis. Drive testing, a long-established approach, involves systematic data collection utilising specialised Radio Frequency (RF) measurement equipment. As exemplified by regulatory bodies like the UK’s Office of Communications (Ofcom) Ofcom (2024a), this often includes sophisticated scanning receivers and calibrated devices capable of concurrently measuring key performance indicators across multiple mobile network operators and technologies (e.g., 2G, 3G, 4G, and 5G) along predefined routes Ofcom (2024b). This method affords a high degree of control over testing parameters and precise geolocation, enabling the acquisition of granular data essential for ensuring regulatory compliance and informing policy-making. The primary strengths of drive testing lie in its precision, the depth of diagnostic information obtainable, and the ability to perform controlled, repeatable measurements. However, its operational costs are substantial due to specialised equipment and personnel, and its spatio-temporal coverage is inherently limited by logistical constraints, often restricting data to major transport corridors.
In contrast, crowdsourced data analysis leverages measurements collected from mobile apps on a multitude of consumer smartphones. This approach offers unparalleled scalability, providing extensive geographic and temporal coverage at a significantly lower operational cost than drive testing. The data typically include network coverage, internet connectivity, device information, and contextual metadata. Crowdsourced apps in particular enable extensive sampling in indoor locations that are often inaccessible to conventional drive testing. Notably, they also record instances of no signal, immediately logging a location when a phone has no coverage. These "no-service" data points highlight complete coverage holes that traditional methods might miss. Analysing these areas may reveal important socioeconomic patterns Koutroumpis and Leiponen (2016) to drive discussions on evidence-based policymaking processes.
@@
Using crowdsourced data, we first contrast a geometric approach for signal coverage analysis with a machine-learning method. We then outline a time-series-based validation strategy and examine the results. Finally, we consider practical engineering applications of improved coverage analysis and their business impact.
2 Methodology
@@
One traditional approach to estimating mobile coverage areas from crowdsourced data is to use computational geometry, for example by constructing a convex hull around all the geolocated points where service was observed. The convex hull provides a simple approximate boundary of the network’s coverage. If a geographic area lies outside this hull, it can be flagged as a potential coverage hole. This method is conceptually straightforward and computationally efficient. Yet, the convex hull method tends to overgeneralise because it cannot represent concavities, internal holes, or irregular coverage boundaries and it is strongly affected by outliers or widely spaced points Neidhardt et al. (2013); Brunello et al. (2023).
@@
In our approach, we model coverage estimation as a one-class classification (novelty-detection) problem: given geolocated measurements that evidence usable service (the "inliers"), we learn the support of that distribution and treat points outside it as likely weak or no coverage. The One-Class SVM (OC-SVM) algorithm Schölkopf et al. (2001) estimates a function $f(x)$ whose sign indicates membership in the learned support. The prediction $f(x)\geq 0$ denotes coverage, whereas $f(x)<0$ indicates a lack of coverage. Unlike convex geometric baselines, the kernelised decision boundary can be highly non-convex, capturing concavities and internal holes driven by terrain, clutter, or shadowing. In practice, we use the Radial Basis Function (RBF) kernel to obtain smooth, locality-aware boundaries that wrap around dense regions of positive evidence without being pulled by isolated outliers. This mirrors common use of one-class SVM in anomaly detection, where the algorithm learns the region occupied by "normal" data and flags deviations as anomalies.
@@
A key advantage of OC-SVM is the soft boundary controlled by the hyperparameter $\nu\in(0,1]$. The $\nu$ parameter simultaneously sets an upper bound on the fraction of training errors (points allowed outside the learned support) and a lower bound on the fraction of support vectors. It controls the trade-off between overfitting (making the boundary too tight, excluding real coverage) and underfitting (making the boundary too loose, including false coverage). A small $\nu$ yields a tighter boundary that may exclude some true coverage points, while a large $\nu$ produces a looser boundary that may include false positives. Thus, $\nu$ governs the model’s sensitivity to outliers and the complexity of the learned support region.
@@
Figure 1: Mobile signal coverage area of a single cell tower. On the left is the convex hull approach (a), which overgeneralises and includes large areas of no coverage. On the right is the OC-SVM approach, which captures concavities and internal holes to better represent the true coverage area. The right figure (b) is a smoothed, non-convex boundary that tightly wraps around the empirical coverage points, calculated using OC-SVM with RBF kernel. Ope (2025)
@@
Meanwhile, the RBF width is governed by $\gamma$ in $K(x,x^{\prime})=\exp(-\gamma\lVert x-x^{\prime}\rVert^{2})$. A large $\gamma$ yields a narrow influence radius, so each support vector only affects classification in its immediate vicinity and the learned decision boundary can become overly sensitive to noise. Conversely, a small $\gamma$ produces a broad influence radius, allowing each support vector to influence a wide region and yielding a smoother, more general boundary that may underfit fine spatial detail. Selecting $\gamma$ therefore sets the granularity of the learned surface.
@@
To select the best hyperparameters, we tune $(\nu,\gamma)$ by temporal cross-validation using held-out time slices. We partition the data into training and validation sets along the temporal dimension (e.g., train on January measurements and validate on February). This simulates real-world deployment, where future coverage must be predicted from historical observations. For each candidate $(\nu,\gamma)$ pair, we train the OC-SVM on the training set and evaluate its performance on the validation set using metrics like recall (true positive rate) and precision (positive predictive value). We select the hyperparameters that yield the best trade-off between recall and precision on the validation set, ensuring robust generalisation to unseen data.
@@
In our approach, we partition the crowdsourced measurement data by signal levels and train a separate OC-SVM boundary for each level. The intuition is that the spatial support of good coverage points will differ (be more conservative) from that of weaker signal and so by modelling each level’s support separately, we can obtain layered (nested) boundaries that demarcate zones of stronger versus weaker coverage. Each OC-SVM is responsible for estimating the region in which that signal level is reliably observed. In deployment, one can interpret a location’s highest predicted signal level by querying which OC-SVM boundary it falls inside.
@@
We compare our OC-SVM approach against the convex hull baseline by evaluating both methods on the same temporal validation splits. We assess their ability to correctly identify known coverage areas (true positives) and avoid falsely predicting coverage in known no-service areas (false positives). Metrics like recall, precision, and F1-score provide a quantitative comparison of their performance. We expect the OC-SVM to outperform the convex hull by better capturing complex, non-convex coverage boundaries and reducing false positives due to its learned, data-driven nature.
@@ 3 Results
We conducted a grid search over $\nu\in\{0.02,0.04,0.06,0.08\}$ and $\gamma\in\{1\times 10^{4},2\times 10^{4},3\times 10^{4},4\times 10^{4}\}$ using time-slice cross-validation. As illustrated in Figure 2, lower values of $\nu$ produced tighter decision boundaries with fewer outliers, but often overfit the training data. In contrast, larger $\nu$ values allowed more points to lie outside the boundary.
Figure 2: Effect of varying hyperparameters $(\nu,\gamma)$ on the shape and fidelity of the OC-SVM boundary. Note the trade-off between smoothness and fragmentation.
@@
We systematically evaluate the performance of OC-SVM boundaries using various hyperparameters against the convex hull baseline. The choice of F1 score is particularly useful in our context because it provides a single metric that balances both precision and recall. This is important because, in coverage area estimation, we want to minimise both false positives (areas incorrectly marked as covered) and false negatives (missed coverage areas). Using the F1 score ensures that our model performs well across both metrics rather than optimising for one at the expense of the other. The F1 score formula is given by:
@@
Category Signal Level (dBm) Convex Hull OC-SVM
@@
We performed cross-validation by partitioning the data into training and validation sets along the temporal dimension. For each candidate pair of hyperparameters $(\nu,\gamma)$, we trained the OC-SVM on the training set and evaluated its performance on the validation set. The convex hull method was applied to the same training data for a fair baseline comparison. Each model was trained for individual cell towers to capture their unique coverage characteristics. We analysed more than 4,000 cell towers across England for our investigation, ensuring a diverse representation of urban and rural environments. The results in Table 1 and Figure 3 are averaged over all cell towers. We observed that the OC-SVM approach consistently outperforms the convex hull approach, particularly in scenarios with variable or poor signal levels where coverage boundaries are more complex. This demonstrates the effectiveness of the OC-SVM in capturing non-convex coverage areas and reducing false positives.
@@ 4 Discussion
This study set out to address the challenge of accurately estimating mobile network coverage areas using crowdsourced data. The primary aim was to evaluate whether the OC-SVM could provide more precise and reliable coverage boundaries compared to a more traditional geometric method using a convex hull. By doing so, we sought to improve the identification of coverage gaps and weak spots, ultimately supporting better network planning and QoE.
The results demonstrate that the OC-SVM approach consistently improves on the accuracy of coverage boundary estimation compared to the convex hull approach. This is particularly significant in urban environments where coverage gaps are often hidden by complex terrain and building structures. The kernelised decision boundary captures complex, non-convex shapes that more accurately reflect real-world coverage, improving precision because the model is less likely to include areas without service and enhancing recall because it better identifies true coverage areas.
Numerous attempts have utilised Machine Learning (ML) techniques to enhance traditional path loss models Zhang et al. (2019); Sousa et al. (2021); Isabona et al. (2022). However, applying an OC-SVM classifier to crowdsourced data for coverage estimation is a novel contribution that extends the use of ML in wireless communications. This approach draws on ML’s strength in handling complex, non-linear relationships, making it well suited to capturing the intricacies of mobile network coverage that traditional methods may overlook. Incorporating crowdsourced data also provides an empirical basis for understanding real-world coverage patterns, rather than relying solely on theoretical models or limited drive-test data that are both expensive and time-consuming to collect. Additionally, the utilisation of crowdsourced measurements allows for a dynamic and up-to-date representation of network coverage, reflecting evolving changes in the environment and user behaviour that static models may miss, such as new building developments, changes in user density, and seasonal variations in the environment.
@@
The improved accuracy of coverage boundary estimation has significant implications for network planning and optimisation. By accurately identifying coverage gaps and weak spots, operators can make informed decisions about where to deploy additional resources to address service gaps, such as new cell sites or signal boosters, and to select appropriate spectrum bandings for both rural and urban areas. This targeted approach can lead to more efficient use of resources and ultimately enhance the QoE for end users.
@@ 5 Conclusion and Future Work
This study proposes the OC-SVM as a method for mobile network coverage analysis using crowdsourced geolocation data. We compared the proposed approach against a geometric convex hull baseline and demonstrated that kernelised OC-SVM can effectively capture complex, non-convex coverage boundaries that better reflect real-world signal propagation, particularly in urban environments with significant clutter, terrain, and shadowing effects. With appropriate tuning, the SVM boundaries achieve a balance between coverage inclusivity and the exclusion of spurious outliers, enabling the estimation of coverage boundaries at various signal strength levels. The results confirm that crowdsourced data, despite inherent heterogeneity and noise, can be transformed into reliable coverage models when combined with suitable machine learning techniques. This offers direct value both to operators seeking to prioritise investment and to regulators monitoring equitable service provision.
@@
Because kernelised OC-SVM natively supports multi-dimensional inputs, the proposed approach can be extended to 3D geolocation data (i.e., longitude, latitude, altitude). This could further enrich coverage analysis, particularly in dense urban environments where verticality plays a key role.
@@
In summary, the proposed method demonstrates that a machine-learning approach can be used to estimate mobile coverage boundaries from empirical crowdsourced data. This opens up new possibilities for network analysis, optimisation, and regulation. With further refinement and extension, this framework has the potential to become a standard tool for mobile network coverage assessment in the era of machine learning.
@@
Each measurement record includes information such as the timestamp, geographic coordinates, radio technologies (2G, 3G, 4G, and 5G), and key performance indicators including received signal strength, throughput, latency, and service availability. These metrics reflect the network’s performance as experienced by users, rather than being inferred from network-side counters. All data are anonymised, aggregated, and processed in compliance with data protection and privacy frameworks before being made available for analysis.
For this study, a synthetically generated dataset based on UK NetPerform was used Feehily et al. (2025). The synthetic data comprise geolocated mobile signal measurements spanning multiple months, capturing a broad range of environmental and temporal conditions. The synthetic dataset underwent standard pre-processing, including coordinate validation, outlier removal, and spatial filtering to exclude duplicate events.